Using National Income Account’s sources of spending and uses of income produced perspectives derive and interpret three (3) sectoral balances.


1
Expert's answer
2021-10-27T08:09:36-0400

National income accounting is a way that government measures the country’s economic activity by having a bookkeeping system, which offers insight into the economic performances.

The primary use of National income accounting is a tool in economic policy by measuring the economic activity of a country including GDP (Gross domestic product)

In GDP, flows are derived from the National Accounting relationship between aggregate spending and income

Y = C + I +G+ (X-M)

Where Y  is GDP (expenditure), C is consumption spending, I  is private investment spending,  G is government spending,  X is exports and M  is imports (so X − M = net exports).

Another perspective on the national income accounting is to note that households can use total income ( Y) for the following uses:

Y = C+ S+ T

Where  S is total saving and T  is total taxation (the other variables are as previously defined).

You then bring the two perspectives together (because they are both just “views” of Y  to write:-

C +S+T = Y= C+I+G+(X-M)

Where C can be dropped in both sides to get

C+T = I + G + (X- M)

Then you can convert this into the following sectoral balances accounting relations, which can be rearranged to get the accounting identity for the three sectoral balances – private domestic, government budget and external:

(S – I) + (T –G) = (X-M)

This implies that deficits at home (private and government) result in current account or trade deficits, and thus borrowing from abroad.

The sectoral balances equation, says that total private savings ( S ) minus private investment ( I) has to equal the public deficit (spending,  G minus taxes,  T) plus net exports (exports  X) minus imports (M)), where net exports represent the net savings of non-residents.